Emily Johnson is a 38 year old mother of 2 who presents to her primary care physician’s office with rapid heartbeat, rapid breathing, dizziness, and fatigue. During initial intake, Emily states that she began the Military Diet after her being disappointed by her weight at her last doctor’s appointment. She weighed 196 pounds at her appointment and since starting the diet last week she now weighs 188 pounds. She has been following the “Military Diet” which consists of half a grapefruit and black coffee for breakfast, 1 slice of toast, ½ cup tuna with coffee for lunch, and 3 ounces of grilled chicken, ½ apple, and 1 cup of green beans with tea for dinner. She has been following this eating pattern for 7 days. She reports that she is not exercising, and prior to this latest diet, she was eating on the go most meals, frequently dining out or driving through fast food restaurants. Her doctor determines that she is severely dehydrated, and recommends that she see you to create a healthy, safe, weight loss program. Emily height is 5’4 inches. • What is missing from this diet plan? Discuss 3 ways Emily could achieve a better nutrient balance. [Write your response here] • According to the Mifflin-St. Jeor equation, what are Emily’s current energy needs? To lose about 1 pound per week, what should her calorie goal be? [Write your response here] • Help the patient set a SMART goal (Specific, Measurable, Attainable, Relevant, and Timely) to help her make a healthy lifestyle change. [Write your response here] • What are the physical activity recommendations you would make for Emily to help her achieve her goal? [Write your response here
Answer #1

Age of Emily is 38years

Weight of Emily is 188lbs

Height of Emily is 5feet 4 inches = 64inches

#. The main nutrient component missing in Emily's diet is carbohydrates because fish ,benas ,chicken all provide protein . Fruits provides vitamins and minerals and coffee stimulates the brain as it contains caffeine.

Three ways to prepare a good diet plan are :-

- include carbohydrates ,protein and fat in proportion in the diet

- decrease coffee intake

- including more pulses and fluids

#. According to Mifflin -St.Jeor equation

Emily's current energy needs = (4.55×weight in lbs)+(15.88×ht.in inches)- (5×age in years) - 161

= (4.55×188)+(15.88×64)-(5×38)-161

= 855.4 + 1016.3- 190- 161

= 855.4+ 826.3- 161

= 855.4 + 665.3

= 1520.7 calories is needed by Emily.

To lose 1pound every week her calorie goal should decrease by 4calories every week means if at present her total calories is 1520.7 ,then to decrease 1pound she should take only 1516.1 calories.

#. SMART goal for Emily is :-

S(specific ) - the goal must have its main focus on the weight loss by modifying the diet and including physical activity .

M(measurable) - the goal or aim should be measurable like here the main focus is weight ,she should weight herself every weekly to evaluate the outcome that is lose in weight and she can use it as a motivation to lose more weight .

A(attainable ) - a goal must set in such a way that is attainable . It should not be something that cannot be achieved . Here the gaol is to attain weight lose which is practical on nature.

R ( relevant ) the goal must be relevant according to the needs of the patients .

T( timely ) means the goal of losing weighting should be attained timely, so that there is no lag or stretch in the outcome.

#. Physical activity recommended for her are :-

- brisk walking

- cycling

- swimming

- yoga

- exercise

- doing household activities than being a sedentary housemaker.